Ben is twice Joes age the sum of their ages equal 57 how old is Joe and how old is Ben

Answer by Maths68(1474)

Let
Present age of Ben = b
Present age of Joe = j

Ben is twice Joes age
b=2j.......(1)
The sum of their ages equal 57
b+j=57.......(2)
Put the value of b from (1) to (2)
2j+j=57
3j=57
3j/3=57/3
j=19
Put the value of j in (1)
b=2j
b=2*19
b=38
Joe is 19 and Ben is 38 years old.